Monday 12/03/12

8:45 A — 7:30 P, 8:45P – 12:40 A

Confocal data

  • Finish O/N confocal run
  • transfer files
  • modify imviewer LSM to automatically detect all .lsm files in a folder and convert them all to TIF / labeled stacks of TIFs automatically.
  • start code to analyze dot distances for Dfd-Tub data.
    • Would be better with clear nuclear outlines.  Should have stained nuclear lamina.
    • Should also take overview images a 40x full field (or 20x?) for better staging.
    • Upgrading Matlab to 2012b.

Fly work

  • water recombineering crosses (need new syringe + ddH2O).
  • set up new cage of MTDs
  • MTD x Su(Z)12s F1s so far no viable eggs.  Should collect F1s and fix some embryos

Cryosection staining / DNA-FISH

  • staining worked well, sections still flat
  • brief SSCT wash, brief PBT wash, block 1hr.
  • 50 min in primary antibody (rb H3K9me2 + m Dm01-Cy3B)
  • 1 hr (not more) wash in PBT (minimize opportunity for sections to detach).
  • incubating in secondary (Abcam anti-rb 750).
  • check on Turnkey, no 405 label obvious.
  • Fails 750 on STORM2 (see below).
  • Restaining O/N rb-488.  (how could the abcam 750 fail?  or maybe the dyes have died?)

Wholemount embryo staining / DNA-FISH

  • DNA FISH O/N program finished, samples at RT.
  • exchanged labeling solution for post hybe, take sample of embryo to check labeling on Turnkey. — labeling looks good.
  • Started 37C rinse in post-hybe 2 (50% post-hybe 1 + 40% formamide in PBT) (~20 min)
  • 20% formamide in PBT rinse also at 37C   (~20 min).
  • Rinsing at RT in PBT.
  • Block, primary antibodies rb H3K9me2 + m-Dm01 (unlabeled) at RT 2+ hrs.
  • Rinse RT
  • secondaries: a-m cy3B, abcam anti-rb 750
  • Embedded embryos in small volume of 12% gelatin.
  • Incubating in 50% sucrose solution O/N (should have saved some embryos to check on confocal).

Probe Making

  • pour PAGE denaturing gel to separate bands for probe.
  • Remember to heat denature first this time!
  • Still only single band.  Maybe denaturing gel did work last time.  Last time band was much tighter.  160 uL of probe with + 170 uL 2x loading buffer to nearly 400 uL sample needed two wells.  not as bright and a little more spread out.
  • Gel crushed and DNA redissolving at 55C in 2 tubes per band O/N.
  •   Will precipitate and test tomorrow.
  • Send Order probe library from LC to Matt for PO.
In situs day 1
  • using YW O/N embryo stock
  • inv/en-DNP, tou-bio, E(pc)-dig
  • dfd-bio, tub-dig
  • use directly labeled Dmo1 to mark nuclear lamina
  • got poly-lysine from hazen (Wenqin, properly stored, never opened, newly recieved)
  • high static light white plastic/styrophome like crystal.  
  • made .5 mg/mL (25 mg in 50 mL) stock solution in PBS pH 8.5 (probably should have used tris).  coated coverslips — dry cloudy.


  • centromere stain switching very well.  Some sections well attached.
  • No trace of 750 (or 405 on turnkey) at all
  • staining O/N at 4C with a-rb 488.   I don’t understand why this doesn’t work.


  • O/N imaging of Dfd/tub
  • took overview of full field 40x (70%-whole embryo) for staging.  Imaging 3 channels at speed 6, 22 positions, 35 heights.  Hopefully should be done by 10A.
This entry was posted in Summaries and tagged , , . Bookmark the permalink.